### v1

Abstract
The common observation of a ‘$1/f^{\alpha}$’ power spectrum with $\alpha<2$ constitutes one of the enduring mysteries of condensed matter physics. Here it is shown that a $1/f^{\alpha}$ power spectrum, with $\alpha=3/2$, can arise when an ensemble of two–level systems is coupled to a heat bath by means of a system of Bosonic quasiparticles. The model considered is the classic model of Faughnan and Strandberg of the phonon bottleneck, and the anomalous relaxation is associated with an approximate non-equilibrium steady state of the phonons maintained by slow spin relaxation. It is shown that a frequency-dependent susceptibility can be defined in the steady state and that the prediction $\alpha=3/2$, or the equivalent stretched exponential relaxation with exponent $\beta=2-\alpha=1/2$, is consistent with existing experimental data. These results give an insight into the origins of anomalous relaxation in condensed matter and the practical behaviour of two-level systems.
Exponential, or Debye, relaxation has a power spectrum that tends to $1/f^{2}$ at high frequency $f$. However many real condensed matter systems – conducting [1, 2], dielectric [3, 4, 5] and magnetic [6, 7, 8, 9, 10, 11] – show power law behavior $\sim 1/f^{\alpha}$ with an exponent $\alpha$ that lies in the range $1\leq\alpha<2$. While a number of particular mechanisms for this behavior have been established (see e.g. Refs. [6, 7, 8, 9, 10] for the magnetic case), in general it remains a mystery as to why $1/f^{\alpha}$ behavior is so widespread. A common explanation [1] is that the power law is constructed from a wide range of exponential timescales, but in physical terms the range of timescales required (many decades) is usually unrealistic. An exception might be where the time scales are associated with an underlying broad range of length scales, as occurs in a critical or fractal system, but such instances cannot account for the majority of observations. This leads one to suspect that there are more fundamental causes at play, arising in non-equilibrium physics and quantum dynamics. In this context, simple and realistic models and experiments that produce well-defined exponents $\alpha$ are of evident interest to signpost the way to fundamental causes.
This paper reconsiders the canonical model of the ‘phonon bottleneck’ due to Faughnan and Strandberg [12]. The model exemplifies a simple non-equilibrium, and essentially quantum mechanical, model that describes an ensemble of two-level systems (spins) connected to a heat bath by an intervening system of Bosonic quasiparticles (phonons). In a certain parameter regime the model describes a non-equilibrium steady state, and it is shown here that the steady state is associated with almost perfect $1/f^{3/2}$ power spectrum. This result, derived below, is illustrated in Fig. 1 where the exact solution for $|\chi^{\prime\prime}|/\omega$ (the ratio of the imaginary part of the magnetic susceptibility to the angular frequency, a measure of the power spectrum [13]) is compared to the approximation :
$\frac{|\chi^{\prime\prime}|}{\omega}=\frac{\chi_{T}}{\sigma^{-1}+\sqrt{2\sigma}\,\omega^{3/2}}$ (1)
(where $\chi_{T}$ is the isothermal susceptibility), and a $1/f^{3/2}$ power spectrum (given $f=\omega/2\pi$) is observed over many decades of frequency.
Figure 1: Exact power spectrum (blue line) compared with Eq. 1 (magenta dashed line). Here $\sigma=10^{5}$ and a power law region with $\alpha=3/2$ extends over $1/\sigma<\omega<\sigma$. The two curves are almost coincident on a log–log scale, illustrating a rare instance where a ‘$1/f^{\alpha}$’ power spectrum can be calculated analytically (here $\alpha=3/2$ and $f=\omega/2\pi$).
The model considers a dilute system of two-state spins that relax by directly emitting or absorbing phonons in a narrow band around the resonant frequency. Owing to ‘quantum freezing’ of phonon modes at low temperatures, the phonon specific heat becomes too small to maintain thermal equilibrium, hence the ‘bottleneck’. The phonons are treated quantum mechanically [14], leading to two coupled equations of motion.
$\frac{dv}{dt}=-v+\sigma u(1-v)$ (2)
$\frac{du}{dt}=a\left[(v-u)-\sigma u(1-v)\right]$ (3)
Here the variable $v$ is defined as
$v=\frac{(N_{i}^{0}-N_{j}^{0})-(N_{i}-N_{j})}{(N_{i}^{0}-N_{j}^{0})}$ (4)
where $N_{i}$ is the number of spins in the excited state, $N_{j}$ is the number of spins in the ground state, the superscripts represent equilibrium values and $N_{i}+N_{j}=N_{i}^{0}+N_{j}^{0}=N$, the number of spins. If $i$ denotes spin up, and $j$ denotes spin down, then the magnetic moment $\mathcal{M}=N_{i}-N_{j}$ and the equilibrium magnetic moment is $\mathcal{M}^{0}=N_{i}^{0}-N_{j}^{0}$. It follows that $v=(\mathcal{M}^{0}-\mathcal{M})/\mathcal{M}^{0},$ a normalised measure of the deviation of magnetic moment, or magnetization, from its equilibrium value. For example, if a field is suddenly applied to an equlibrium unmagnetized system, then $v$ will evolve from $v=1$ to $v=0$ as the new equilibrium is established at $\mathcal{M}=\mathcal{M}^{0}$.
The variable $u$ is defined as
$u=\frac{1}{\sigma}\frac{p-p^{0}}{p^{0}+\frac{1}{2}}$ (5)
where the $p$’s are mean phonon occupation numbers at the frequency corresponding to the level splitting. Finally the two parameters are defined as follows. The parameter $a=\tau_{\rm s}/\tau_{\rm p}$ where $\tau_{\rm s}$ is a spin relaxation time and $\tau_{p}$ is a phonon relaxation time. The parameter $\sigma$ is defined by
$\sigma=\frac{c_{s}/\tau_{s}}{c_{p}/\tau_{p}}.$ (6)
where the $c$’s are specific heats. A large value of $\sigma$ indicates a relatively small phonon specific heat: hence it is a measure of the strength of the ‘phonon bottleneck’. Note that, in this treatment, $\sigma$ and $a$ are dimensionless, so the time $t$ and angular frequency $\omega$ are also dimensionless, defined with respect to some reference time scale.
The situation of interest arises where $\sigma,a\gg 1$. The phonons are strongly bottlenecked but the spins relax relatively slowly, continually maintaining a far from equilibrium phonon population. This leads to an approximate steady state where the rate of change of the phonon population can be neglected. Thus $|du/dt|\ll a$, and the right hand side of Eq. 3 may be set to zero. This reduces Eqs. 2, 3 to a single equation depending only on $\sigma$:
$\frac{dv}{dt}=\frac{-v}{1+\sigma(1-v)}=-u.$ (7)
At short times $v\rightarrow 1$, but for sufficiently large $\sigma$, there appears to be a region at $1-v\gg 1/\sigma$ where this equation approximates:
$v\approx 1-\sqrt{2t/\sigma}\approx e^{-\sqrt{2t/\sigma}}.$ (8)
Correspondingly, the Fourier transform of this function at high frequency should approximate
$v(\omega)\approx\frac{1}{\sqrt{2\sigma}\omega^{3/2}},$ (9)
suggesting an equation analagous to Eq. 1 for $v(\omega)$.
These ideas are born out by numerical solution of the coupled equations. Using the example of $a=10^{2},\sigma=10^{5}$, Fig. 2 (upper) illustrates the solution for $v(t)$ compared to $1-\sqrt{2t/\sigma}$ and to $e^{-\sqrt{2t/\sigma}}$, which are very close approximations at small times, and the (exact) inverse Fourier transform [15] of the right hand side of Eq. 1, which captures the whole curve. Fig. 2 (lower) is the corresponding numerical Fourier transform on a log-log scale, confirming a slope very close to 3/23/23 / 2. Fig. 3 illustrates the corresponding $u(t)$ compared to $\frac{d}{dt}e^{-\sqrt{2t/\sigma}}$ (see Eqs. 7 and 8) which is a close approximation at short time after transient effects have died out.
Figure 2: Relaxation of the magnetization difference: numerical solution to Eqs. 2, 3, for $a=10^{2},\sigma=10^{5}$. (a) the function $v(t)$ (blue line) compared to $1-\sqrt{2t/\sigma}$ (gray dotted line), $e^{-\sqrt{2t/\sigma}}$ (gray dashed line) and to the (exact) inverse Fourier transform of the right hand side of Eq. 1 (magenta dashed line). (b) Fast Fourier Transform of $v(t)$ (blue points) compared to a line of slope 3/23/23 / 2 (orange line) on the log-log graph.
Figure 3: Relaxation of the phonon population: $u(t)$ numerically calculated in the example of Fig.2 (red dashed line) compared to the prediction $w\approx\frac{d}{dt}e^{-\sqrt{2t/\sigma}}$ (see Eq. 7). There is a close agreement once transient effects have died out at very short time.
Exact solutions to Eqs. 2, 3, 7, may be explored by rewriting the equations as a function of the magnetization $M=\mathcal{M}/V$, where $V$ is the volume. Eqs. 2,3 become
$\frac{dM}{dt}=(M^{0}-M)-\sigma uM$ (10)
$\frac{du}{dt}=a\left(1-\frac{M}{M^{0}}-u-\sigma u\frac{M}{M^{0}}\right)$ (11)
and Eq. 7 becomes:
$M^{0}-M=\frac{dM}{dt}\left(1+\frac{\sigma M}{M^{0}}\right).$ (12)
In the limit that Eq. 12 applies, now consider that the position of equilibrium is set by an applied field $H=H_{1}+he^{i\omega t}$. We have $M_{0}=\chi_{T}H$ where $\chi_{T}$ is the equilibrium isothermal susceptibility, and writing $M=M_{1}+m(\omega)e^{i\omega t}$ we may define a complex susceptibility $\chi(\omega)=\partial m/\partial h$. After differentiating Eq. 12 with respect to $h$, Eq. 12 becomes:
$\chi_{T}-\chi=i\omega\chi\left(1+\sigma\frac{\chi}{\chi_{T}}\right),$ (13)
from which the real and imaginary parts of $\chi$ can be separated:
$\chi^{\prime}=\frac{-\omega\chi_{T}+\chi_{T}(1+(2+16\sigma(1+\sigma))\omega^{2}+\omega^{4})^{1/4}\sin\phi}{2\sigma w}$ (14)
$\chi^{\prime\prime}=\frac{\chi_{T}-\chi_{T}(1+(2+16\sigma(1+\sigma))\omega^{2}+\omega^{4})^{1/4}\cos\phi}{2\sigma w}$ (15)
where $\phi=\frac{1}{2}\arg[1+i(2+4\sigma+i\omega)w]$ and $\cos\phi,\sin\phi\rightarrow\frac{1}{\sqrt{2}}$ for large $\sigma$.
According to Eq. 15, for large $\sigma$, the function $|\chi^{\prime\prime}|/\omega$ tends to $\chi_{T}\sigma$ as $\omega\rightarrow 0$ and tends to
$\frac{|\chi^{\prime\prime}|}{\omega}\rightarrow\frac{\chi_{T}}{\sqrt{2\sigma}\,\omega^{3/2}},$ (16)
for large $\omega<\sigma$ (cf Eq. 1, Eq. 9). This justifies the close approximation of the exact form by Eq. 1, as displayed in Fig. 1. However, the high order terms in $\omega$ in Eq. 15 ensure that the true asymptotic behavior of $|\chi^{\prime\prime}|/\omega$ as $\omega\rightarrow\infty$ is $1/\omega^{2}$, with a crossover to this form at $\omega\sim\sigma$. Thus the anomalous ‘$\alpha=3/2$’ scaling applies to the range $1/\sigma<\omega<\sigma$ and only the divergence of $\sigma$ would extend its range to infinity. Nevertheless, as illustrated in Fig. 4, for finite $\sigma$, the crossover is particularly sharp and has negligible effect on the power law scaling at smaller $\omega$. Were such behavior encountered in experiment, it would surely be interpreted as an asymptotic $1/\omega^{3/2}$ power law as one would have to access unfeasibly small values of $|\chi^{\prime\prime}|/\omega$ to see the crossover.
A similar analysis of the coupled equations shows that the existence of a frequency-dependent susceptibility $\chi(\omega)$ strictly implies either that $dw/dt=0$ (the case already discussed) or $a=0$, where Debye relaxation is recovered. Thus the two equations become:
$u=\frac{\chi_{T}-\chi-i\omega\chi}{\sigma\chi}$ (17)
$\frac{du}{dt}+u\left(1+\sigma\frac{\chi}{\chi_{T}}\right)=a\left(1-\frac{\chi}{\chi_{T}}\right)$ (18)
and a solution of the lower equation with $u=0$ at $t=0$ gives:
$u=\frac{\chi_{T}-\chi}{\chi_{T}+\sigma\chi}\left(1-e^{-at\left(1+\sigma\chi/\chi_{T}\right)}\right)$ (19)
which is consistent with Eq 17 (whose right hand side does not involve time) only in the limit $at\rightarrow\infty$. However, at $\omega\sim 1/t\ll a$, corrections to the limiting solution become exponentially small and the analytic form Eqs. 1, 15 should represent a close approximation to the solution of the coupled equations, as already observed (Fig. 2).
Figure 4: The sharp crossover between $\alpha=3/2$ to $\alpha=2$ scaling at $\omega\sim\sigma$, with $\sigma=10^{5}$ in this example. Here the blue line is $|\chi^{\prime\prime}|/\omega$ with $\chi^{\prime\prime}$ given by the exact expression, Eq. 15, The magenta dashed line is Eq.1 with $\alpha=3/2$, and the grey dashed line is $1/\omega^{2}$ with $\alpha=2$.
The limits $a\rightarrow 0$ ($u=0$) and $a\rightarrow\infty$ ($u$ finite) thus constitute two steady states where a frequency dependent susceptibility can be defined. In the former case the phonons are in equilibrium with the bath and have the Bose distribution $p^{0}=1/(e^{h\nu/kT}-1)$, while in the latter case the phonons have a non-equilibrium distribution.
Stretched exponential relaxation (Eq. 8) does not seem to have been tested experimentally for the phonon bottleneck in dilute spin systems, despite non-exponential relaxation being widely expected (and occasionally observed) following Ref. [12]. The many studies (see Ref. [16] for early work) have generally focused on temperature dependence, rather than time or frequency dependence, but there are at least two studies, by Standley and Wright [17], and by Roinel et al. [18], that tested Eq. 7 in the form of its solution for $t$:
$t=-\sigma\left(1-v\right)-(1+\sigma)\ln\left(v\right).$ (20)
Both studies reported excellent agreement with the experimental data. It was, of course, necessary to introduce units of time, by writing $t=\mathcal{W}t^{\prime}$ where $t^{\prime}$ is the dimensional time. Standley and Wright used esr to study the magnetic relaxation of a an organic copper complex (copper dipyrromethene), following a microwave pulse, fitting their data at 2.45 K (for example) with $\mathcal{W}=1.06/{\rm ms}$, $\sigma=1.57$. Roinel et al. used nmr to measure the relaxation of the polarisation (magnetization) of 169Tm nuclear spins in TmPO4 at $T=50$ mK, $H_{1}=1$ T, fitting their data with $\mathcal{W}=1/(26~{\rm hr})$ and $\sigma=17$.
Figure 5: Comparison with of the theory with experimental data on a reduced timescale $t_{0}$ defined in the text. The lower figure is an enlargement of the upper one for short times. The black points are the data of Ref. [17], and the orange points are the data of Ref. [18]. The blue lines are the exact solution, Eq. 21 for (top to bottom) $\sigma=1.57$ [17], $\sigma=17$ [18] and $\sigma=500$, a large value for comparison. The magenta dashed line is the exact Fourier transform of Eq. 1.
Eq. 20 is exact and equivalent to the following exact solution for $v(t)$:
$v(t)=-\frac{1}{1+\sigma}W\left(-\frac{\sigma\left(e^{-\sigma-t}\right)^{\frac{1}{1+\sigma}}}{1+\sigma}\right)$ (21)
where $W$ is the Lambert W-function. This equation shows that curves of different $\sigma$ can be compared on a scale of reduced time $t_{0}=t/(1+\sigma)$. For the inverse Fourier transform of Eq. 1 the corresponding transformation is $t_{0}=t/\sigma$ [15], but the difference is immaterial given that the latter is a large $\sigma$ approximation.
Rescaling accordingly, the data are replotted in Fig. 5 and compared with the theory. It can be seen that the value of $\sigma=17$ (Roinel et al’s data) is already approaching the large $\sigma$ limit. This is to be expected as, following the earlier arguments, $\sigma=17$ should give over two decades of $1/\omega^{3/2}$ scaling, whereas for $\sigma=1.57$ (Standley and Wright’s data), the corresponding range would be negligible. It would be interesting to find further examples of such behavior among real systems, particularly with much larger values of $\sigma$.
The value of $\alpha=3/2$ found here, equivalent to stretched exponential ($e^{-t^{\beta}}$) relaxation with $\beta=2-\alpha=1/2$, may be compared to the values found in other magnetic systems. Ising spin glasses [6] display values of $\beta\approx 0.25-0.3$. In single molecule magnets, a value of $\beta=1/2$ has been observed at millikelvin temperatures at resonant tunneling points and interpreted in terms of fluctuating random fields [8, 9]. In spin ice, $\alpha$ evolves from 222 down to about 1.5 at subkelvin temperatures [11], where the value of roughly 1.5 has been associated with the diffusion of emergent magnetic monopoles in a spatially fractal environment [10, 20]. It is perhaps unsurprising that that the few-body systems (paramagnets, single molecule magnets) seem to exhibit exponents that are rational fractions, while the many body systems (spin glasses, spin ice) exhibit more general exponent values.
It is finally worth noting some general insights afforded by the present results. Exponential relaxation indicates that the rate of some relaxing quantity is proportional to its displacement from equilibrium. In the present case $dv/dt$ is instead linear with $u$, the displacement of the phonon equilibrium (Eq. 7), but the latter is a highly nonlinear function of $v$, and this is the origin of the approximate $e^{-\sqrt{t}}$ relaxation. A situation that is mathematically (though not physically) similar arises in the case of $e^{-\sqrt{t}}$ relaxation in an ordered two-dimensional Ising model where the radius of relaxing circular domains (droplets) plays a similar role to $u$ in the phonon case [21]. In physical terms, the example of the phonon bottleneck shows how just two relaxation processes (both in themselves exponential) may generate a $1/f^{\alpha}$ power spectrum or $e^{-t^{\beta}}$ relaxation, when sufficiently strongly coupled. The anomalous $\alpha,\beta$ are associated with an emergent, non-equilibrium steady state, where a frequency-dependent linear response function can be defined. The phonon bottleneck thus illustrates how anomalous relaxation can emerge in a basic, but realistic, model of a far from equilibrium, and essentially quantum mechanical, steady state. Given its simplicity, the model might lend itself to generalisations, and given current interest in two-level systems, either as qubits [22] or as sources of noise [23], it might find contemporary relevance.
Acknowledgements.
It is a pleasure to thank P. Holdsworth and S. Giblin for useful comments.